Find the vertex and the axis of symmetry and graph the function
f(x)=x^2-9x

Answer by TimothyLamb(4379)

f(x) = x^2 - 9x
---
the above quadratic equation is in standard form, with a=1, b=-9, and c=0
---
to solve the quadratic equation, plug this:
1 -9 0
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
x-intercepts:
the roots of the quadratic are real
9
0
( 9, 0 )
( 0, 0 )
---
y-intercept:
f(0) = 0^2 - 9(0) = 0
---
vertex min or max
vertex is a minimum at ( 4.5, -20.25 )
---
axis of symmetry:
x = 4.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php